OpenStudy (anonymous):

lim of x as sin(3x)/(2x) approaches 0

OpenStudy (agreene):

try graphing it.

myininaya (myininaya):

do you know how to evaluate \[\lim_{u \rightarrow 0}\frac{\sin(u)}{u}?\]

myininaya (myininaya):

This limit should have been already introduced to you by the squeeze theorem (or at least this is the first way I learned what the above limit is) \[\lim_{u \rightarrow 0}\frac{\sin(u)}{u}=1\] commit this limit to memory it will be useful let's take this limit and see if i can give you a hint on how to do your problem since there is a 3x inside that sin let's let u equal 3x and if u goes to 0 then 3x goes to 0 since u=3x and since 3 doesn't go to 0 then the x must go to 0 so anyways \[\lim_{x \rightarrow 0}\frac{\sin(3x)}{3x}=1 \] try to use this limit here for your problem you may find it necessary to multiply by a 1 to do so